OpenStudy (anonymous):

Solve this equation for x. Round your answer to the nearest hundredth. 6=In(x+4)

OpenStudy (anonymous):

6=ln(x+4).....this is base e raised to power 6=(x+4) have to use calculator for e^6=403.42=(x+4) so x=403.42-4=3.99.42

OpenStudy (anonymous):

3.99.42 is the answer?

OpenStudy (anonymous):

isn't it 400?

OpenStudy (anonymous):

I think I read that wrong

OpenStudy (anonymous):

i have an extra dot in there 399.4 you have to give to nearest 1/100

OpenStudy (anonymous):

so is it 399.43

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you! can you help me with another one?

OpenStudy (anonymous):

Solve this equation for x. Round your answer to the nearest hundredth. 0.52=log x @jamaica25

OpenStudy (anonymous):

oops they went offline. @Austin6i6 can you help me with this?

OpenStudy (anonymous):

sure is it log base 10 ?

OpenStudy (anonymous):

I think so..

OpenStudy (anonymous):

mathway said the answer is 3.311

OpenStudy (anonymous):

yeah thats it

OpenStudy (anonymous):

if you raise both sides by base 10 you can cancel the log

OpenStudy (anonymous):

ok so 3.311 and that's rounded to the nearest hundredth?

OpenStudy (anonymous):

\[10^{.52} = 10^{logx}\]

OpenStudy (anonymous):

no that is roundend to the nearest thousands

OpenStudy (anonymous):

.01 = 1/100

OpenStudy (anonymous):

is it 3.31 then?

OpenStudy (anonymous):

yes

OpenStudy (anonymous):

thank you